User name M Jones
Log entry time 08:12:31 on October27,2002
Entry number 89111
keyword=owl shift summary
Finishing run 2769 last run of Q3D_e40 kinematics.
Run 2770 is scint efficiency run
Run 2771 is Q3D_efN
Run 2772 is Q3D_f00 carbon optics target
Run 2773 is Q3D_f00 LD2 prescale for t4 wrong.
Run 2774 is Q3D_f00 RA coll=open LA coll=open
Run 2775 is Q3D_acc RA coll=open LA coll=6msr
Run 2776 is Q3D_acc RA coll=6 msr LA coll=6msr
put in LH2 4cm
Run 2777 is Q3H_acc RA coll=6 msr LA coll=6msr
Run 2778 is Q3H_acc RA coll=open LA coll=6msr
Run 2779 is Q3H_acc RA coll=open LA coll=open
Run 2780 is Q3H_acc RA coll=open LA coll=sieve
Problems moving to 15cm. Had to reboot IOC.
Put in LH2 15cm. Got beam and the target warmed up.
Realized that LH2 target was still in 'sleep' mode.
This means that 4cm runs had fan at 20Hz instead of 60Hz.
Should be redone.
Run 2781 is Q3H_acc RA coll=open LA coll=sieve (short run)
Run 2782 is Q3H_acc RA coll=open LA coll=sieve
Angle camera for left arm is frozen. Need access to enable.
OTR 12 is out. MCC crew doesn't know how to set it up.